2. Suppose g varies inversely as w, and when w is 8, g is 37. What is g when w is 5.2?
3. The illumination from a light source varies inversely as the square of its distance from the light source.
a. Explain what this means specifically.
b. What is the change in illumination of a lamp you moved from 10 to 15 inches from a table?

Answer by malakumar_kos@yahoo.com(315)


2. Suppose g varies inversely as w, and when w is 8, g is 37. What is g when w is 5.2?
3. The illumination from a light source varies inversely as the square of its distance from the light source.
a. Explain what this means specifically.
b. What is the change in illumination of a lamp you moved from 10 to 15 inches from a table?
2) g varies inversely as w i.e, g = k/w where k is a proportionality
constant.
g = k/w
when w = 8 , g = 37 or k = g*w = 8*37 = 296
when w = 5.2 g =? or g = k/w = 296/5.2 = 2960/52 = 56.92

3)a) As the distance of the light source increases illumination decreases.
b)When distance d = 10 inch i = k/100 or k = 100*i
When D = 15 inch I = k/15*15 = k/225
I = 100i/225 = (4/9)i
Hence illumination at 15 inch will be 4/9 times illumination at 10 inch.